OpenStudy (anonymous):

Find the equation of the line tangent to the graph of f (x) = sqrt(x + 1) at the point (3,2).

OpenStudy (amistre64):

which format? slope intercept of point slope form?

OpenStudy (amistre64):

not that format is important; but rather: the slope (m) = f'(3)

OpenStudy (anonymous):

Either one is fine.

OpenStudy (amistre64):

sqrt(x+1) derives to: 1/2sqrt(x+1) ; and i dont see any chain rule trickery so that should be good

OpenStudy (amistre64):

f'(3) = 1/2sqrt(4) = 1/4 right?

OpenStudy (amistre64):

y-Py = m(x-Px) ; where m=f'(3) and P(3,2)